If (a,b) E AxA, where A is a group, then o((a,b))=o(a)o(b)


1
Expert's answer
2021-05-07T13:27:01-0400

Let "A" be a group. If "o(a)=o(b)=2", then "(a,b)(a,b)=(a^2,b^2)=(1_A,1_A)", and hence "o((a,b))=2". It follows that in general case it is not true that "o((a,b))=o(a)o(b)."